A man standing between 2 cliffs hears the first echo of a sound after 2 seconds and second echo 2 seconds after the initial sound. If the speed of sound be 330m/s, calculate the distance between the two cliffs.
Options:
A) 1650m.    B) 990m.    C) 825m.    D) 660m

Grade:9

1 Answers

Arun
5 years ago

t1=2s ,t2=4s , v=330m/s d=?

step1:find distance b/w a man and cliff one

d1=vt1/2=330x2/2=330m

step2:find distance b/w a man and cliff two

d2=vt2/2=330x4/2=660m

step3:find distance b/w cliffs

d=d1+d2=330m + 660m= 990m

Therefore,the distance b/w cliffs is 990m.